AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 278     Entry time: Fri Feb 25 06:28:12 2022
Author: George, Esther 
Subject: Current Status 

07:30
Some signals on the beam monitor, seems they are optimising it.

System Check OK
Temps OK
White Rabbit fail diff: 24
Leakage Current: 3.48uA
Pressures OK
Data rate ~1Mb/s

Attachment 1: 2022-02-25_07-33-44exRate.png  163 kB  | Hide | Hide all
2022-02-25_07-33-44exRate.png
Attachment 2: 2022-02-25_07-35-11xyTotal.png  203 kB  | Hide | Hide all
2022-02-25_07-35-11xyTotal.png
ELOG V3.1.4-unknown